Carol, Chris, Craig, Gordon
These were accepted.
Report from Computing Executive Group
People should now be aware of what projects they'll be working on for T1
Reports from units.
As promised at the last meeting, photos can now be found on the MPU page.
Please report any missing tools or suggestions for additional tools to Ian.
It was agreed that we should switch to switched power bars rather than unswitched.
This doesn't have PXE boot so MPU will arrange for a copy of the latest cd image to be available.
IMPORTANT:Please ensure that all servers are rebooted as soon as possible.
RAT to investigate condor usage. Suspect that there are only around a dozen users. Possibility of pointing them in the direction of ECDF.
The source of the problem is still not confirmed but it was possibly fibre-channel related. The plan is now to incrementally bring back to a similar state for more investigation. It was agreed that it was strange to see the problem over both filesystems and over a reboot. It was noted that it is possible to exclude any binary from prelink and this could be implemented on the ssh servers. We might exclude prelink altogether from sl6.
We have only just taken on this role and plan to ensure that our involvement is limited to technical issues rather than administration.
Topics for discussion
The outcome of the discussion was that this should be taken to CEG for further consideration. It was agreed that there are 2 main strands to the problem i.e. the forensic investigation vs. how to get the service back. The main suggestions/comments included:
Graham has sent MPU code to allow extended use of rfe on lcfg files e.g. searching and deleting.
The next meeting will be on Wednesday 9th March in room 4.31 chaired by Craig.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh